OpenStudy (anonymous):

factor HOW DO YOU FACTOR TRINOMIALS?? any formulas????? 4x² - 5x - 21

OpenStudy (anonymous):

omg so many peolpe here XD...

OpenStudy (zepp):

We have constants 4, -5 and -21 What you are looking for is what numbers (2) summed, get -5, and multiplifed, -21*4.

OpenStudy (anonymous):

hm.. only mulitples of 21 are 7 and 3, but how would they add up to -5?

OpenStudy (anonymous):

here is solution...

OpenStudy (anonymous):

uhm, how did you exactly get the steps? like what did you do to get to the next step? I don't know how you got the next numbers...

OpenStudy (zepp):

-21*4 = -84 right?

OpenStudy (zepp):

So what numbers multiplied gives -84? I found 6*-14, 7*-12 and some others.

OpenStudy (zepp):

Then 7+-12 = 7-12 = -5, so our two numbers is 7 and -12 Now you can rewrite your trinomials into this form: \(4x^2 +7x -12x -21\)

OpenStudy (zepp):

and by factoring out some stuffs out \(4x^2-12x+7x-21\) \(4x(x-3)+7(x-3)\) \((4x+7)(x-3)\)

OpenStudy (anonymous):

check my attachment you get answer...

OpenStudy (anonymous):

ohhh I see now. kk thanks guys

OpenStudy (anonymous):

welcome